* For the function f(x) = (3x + 5 if x 20 6 if x < 0. find f(-10)
Answer #1

we have f(x) = 32 +5 {275 it azo if ako (01-f puif ie If 280, then f(x) = 3x+5 aco then f(x) = 6 and PI since -1020, So f(-o)